OpenStudy (janu16):

What is the equivalent of pi over 4 radians in degrees?

OpenStudy (janu16):

@jim_thompson5910

Nnesha (nnesha):

\[\huge\rm D = r \cdot \frac{180}{\pi}\] multiply radians by 180 degree over pi D=degree r= radians

OpenStudy (janu16):

what is r

Nnesha (nnesha):

Radian

OpenStudy (janu16):

4?

Nnesha (nnesha):

What is the equivalent of `pi over 4 radians` in degrees?

OpenStudy (janu16):

i dont know that part

Nnesha (nnesha):

the question is convert `pi over 4 radian` to degree what is r (radian)?

OpenStudy (janu16):

its 45 degrees

Nnesha (nnesha):

r= pi/4 \[\large\rm D = \frac{\pi}{4} \cdot \frac{180}{\pi}\]

OpenStudy (janu16):

ik the answer is 45. i figured it out